OpenStudy (anonymous):

Subtract the following polynomials, then place the answer in the proper location on the grid.

OpenStudy (alexandervonhumboldt2):

where are they?

OpenStudy (alexandervonhumboldt2):

3.1x-4.3x=-1.2x 2.8+1.2=4 -1.2x+4